The density of butane, C4H10C4H10, is 2.49×10−3 g/mL2.49×10−3 g/mL. Calculate the number of moles of butane in a 95.5 mL95.5 mL sample.

moles of butane:
Answer #1

We know that density is the ratio of given mass of substance to the volume of substance i.e.

d = Given mass / Volume

Putting the values from the question

2.49×10^-3g/ml = w / 95.5ml

w= 2.49×10^-3 × 95.5 g

w = 0.238g

Now given mass is 0.238g. Number of moles is equals to the ratio of given mass to the molar mass and molar mass of butane is 58.12, so number of moles is given by

n = 0.238 / 58.12

n = 0.004

Number of moles is 0.004